OpenStudy (anonymous):

. Solve, using linear combination. 4x + y = 5 3x + y = 3

OpenStudy (anonymous):

4x + y = 5 3x + y = 3 Isolate one of the variables in one of the equations. I'm going to do the first one for y. Subtract 4x from each side. You're left with y=5-4x Substitute 5-4x in the place of y in the second equation. 3x+(5-4x)=3 Combine like terms. -x+5=3 3-5=-2 -2/-1=2 x=2 Check: 3(2)+(5-4(2))=3 6+5-8=3 11-8=3 3=3 So now that you know what x is, you can solve for y. You already know that y=5-4x y=5-4(2) y=5-8 y=-3 Now check in your equations. x should be 2, y should be -3. 4x + y = 5 4(2)+(-3)=5 8-3=5, Check 3x + y = 3 3(2)+(-3)=3 6-3=3 , Check

OpenStudy (anonymous):

Ah I just saw where you asked for it to be solved via linear combination. Idk what the method I used is called and I'm not sure how it differs from linear combination.

OpenStudy (anonymous):

Well thank you for all of that!! but the answer options are: (0, –3) (0, 5) (2, –3) (–3, 7)

OpenStudy (anonymous):

?? :)

OpenStudy (anonymous):

Okay, then it's in (x,y) format. I gave you the answers. x=2, y=-3. So (2,-3) is your answer.